OpenStudy (anonymous):

Find the area of a rectangle with length of 6 inches and width of 2 feet if anyone can get this right medals will come ur way

OpenStudy (anonymous):

1 square foot

OpenStudy (anonymous):

it would be 144 square inches or 1 square foot

OpenStudy (anonymous):

clumsy fingers :(

OpenStudy (anonymous):

2ft= 24 inch A= 6*24 A=144 inches^2 6 inches 1/2 ft A = (1/2)*2=1ft Area; 1ft or 24 inch

OpenStudy (anonymous):

My mistake. 144" = 1 square foot

jhonyy9 (jhonyy9):

you need changing the masure from inch in feet or from inch in feet and wfter you have maked you need just multiplie the length with width and you will can getting the area of this rectangle